# An Event Structure-aware Generative Model for Biomedical Event Extraction

### v2

@@ Abstract
Biomedical Event Extraction (BEE) is a challenging task that involves modeling complex relationships between fine-grained entities in biomedical text. Most existing BEE models rely on classification methods that ignore label semantics and argument dependencies in the data. Although generative models that use prompts are increasingly being used for event extraction, they face two main challenges: creating effective prompts for the biomedical domain and dealing with events with complex structures in the text. To address these limitations, we propose GenBEE, a generative model enhanced with structure-aware prefixes for biomedical event extraction. GenBEE constructs event prompts that leverage knowledge distilled from large language models (LLMs), thereby incorporating both label semantics and argument dependency relationships. Additionally, GenBEE introduces a structural prefix learning module that generates structure-aware prefixes with structural prompts, enriching the generation process with structural features. Extensive experiments on three benchmark datasets demonstrate the effectiveness of GenBEE and it achieves state-of-the-art performance on the MLEE and GE11 datasets. Moreover, our analysis shows that the structural prefixes effectively bridge the gap between structural prompts and the representation space of generative models, enabling better integration of event structural information.
@@ I Introduction
Biomedical Event Extraction (BEE) is a challenging task that involves identifying molecular events from natural language text, where an event typically includes certain biomedical entities, such as genes, proteins, or cellular components [1]. BEE is typically divided into two subtasks, namely biomedical event trigger detection and biomedical event argument extraction. Biomedicalevent trigger detection aims to identify trigger words and classify them by event types that represent the presence of biomedical events. On the other hand, biomedical event argument extraction aims to identify the arguments and associate them with specific roles in biomedical events, correlating them with the corresponding event triggers. Moreover, BEE can provide invaluable information to facilitate the curation of knowledge bases and biomolecular pathways [2].
Traditionally, BEE has been formulated as a classification problem [3, 4]. However, these classification-based methods often struggle with encoding label semantics and other forms of weak supervision information [5, 6, 7]. Moreover, many current methods extract events in a pipeline manner, where trigger detection and argument extraction are treated as separate phases, which hinders the learning of shared knowledge between subtasks [8]. Recently, generation-based models have been proposed for event extraction (EE) on general domains [9, 10]. These models typically cast the event extraction task as a sequence generation problem using encoder-decoder-based pre-trained language models (PLMs) [11, 12] to output conditional generation sequences. These generation-based models benefit from natural language prompts to learn additional semantic information [13]. Apart from employing natural language prompts to enhance semantic learning from the given context, some recent works such as GTEE [14] and AMPERE [15] have also proposed injecting continuous prompts [16, 17] into their frameworks for better contextual representation learning for event argument extraction.
@@
However, most current generation-based EE models [14, 13, 15] have difficulties when dealing with events with complex structures, such as nested events. Compared to flat event structures, texts in the biomedical domain often contain more nested event structures [18]. For example, in benchmark datasets like MLEE [19] and GE11 [20], the nested event structures constitute approximately 25% and 28% of all events in the datasets, respectively. Figure 1 shows an example of two nested events identified from the given text. The trigger word “promoters” of the Positive Regulation event (E2) serves as the argument role Site for the trigger word “bind”, which triggers the Binding event (E1). Furthermore, “TCF-1 alpha” serves as the argument role Theme for both events (E1 and E2). In addition, a trigger word may also trigger different events in a sentence, depending on its event type or arguments. Events with such structures are called overlapping events. The rest of the events are regarded as flat events. Identifying these complex structures is non-trivial, as it can reveal the underlying relationships between events. For example, a Site2 argument in a Binding event is likely to be a trigger for a Positive Regulation event.
For generation-based EE models, an intuitive way to incorporate structural information through prompt engineering [21] is by adding natural language prompts that describe complex structural information as part of the input to the generative PLM’s encoder. This helps elicit the rich knowledge distributed in the PLM. However, Fei et al. [22] pointed out that the mainstay contextual representations in a generative language model’s encoder tend to weaken other high-level information such as structural information, thereby reducing the effects of such information at the decoder. Moreover, it is also unclear whether generative PLMs, such as BART, contain sufficient domain knowledge to represent complex biomedical events. Therefore, incorporating structural information effectively into generative BEE models is challenging.
To address the above issues, we propose an event structure-aware Generative Biomedical Event Extraction model (GenBEE), which can capture complex event structures in biomedical texts. First, GenBEE constructs type description prompts to integrate the semantics of type labels. Next, it distills knowledge from large language models (LLMs), such as GPT-4 [23] to create event template prompts, which incorporate information such as dependencies among argument roles, to enhance contextual representation learning. Then, GenBEE introduces a structural prefix learning module that generates structure-aware prefixes, enriched with structural prompts to infuse the structural features.
@@
  • We propose GenBEE, a structure-aware generative model for biomedical event extraction. Our proposed end-to-end model can distill knowledge from the LLM to incorporate label semantics, dependency relationships among arguments, and shared knowledge between subtasks.
  • In the GenBEE model, we introduce a structural prefix learning module to guide the generation process. This module constructs the structure-aware prefix to link continuous prompts with the representation space of generative models. It is worth noting that this plug-and-play module can be added to any biomedical event extraction framework.
@@ II-A Biomedical Event Extraction
Many current methods for biomedical event extraction adopt a “pipeline” strategy, where trigger detection and argument extraction are treated as separate phases. However, “pipeline” methods may result in the propagation of errors. To address this issue, Trieu et al. [8] introduced an end-to-end method called DeepEventMine, which simultaneously identifies triggers and allocates roles to entities, thereby mitigating error propagation. However, DeepEventMine requires the full annotation of all entities, which may not always be accessible in datasets. Ramponi et al. [6] considered biomedical event extraction as a sequence labeling task that involves joint training models for trigger detection and argument extraction via multi-task learning. Aside from these classification-based methods, some recent works have formulated event extraction as a question-answering task [24, 25]. This new paradigm transforms traditional classification methods into multiple questioning rounds, producing a natural language answer about a trigger or an argument in each round. Current QA-based event extraction methods are primarily focused on formulating separate questions for different events and argument types. However, these QA methods lack the capability on capturing dependency information between arguments, which could be helpful for assigning argument types.
@@ II-B Generative Event Extraction
Recently, another emerging line of work that treats event extraction as a conditional generation problem [26, 27, 28, 13] has evolved. The development of generation-based event extraction models begins with the reformulation of event extraction problems as a generation task. For example, TANL [26] treated event extraction as translation tasks with augmented labels using a variety of brackets and vertical bar symbols for representation, and tuned the model to predict these augmented labels. TempGen [27] proposed using templates for constructing role-filler tasks for entity extraction, and generating outputs that fill role entities into non-natural templated sequences. To adapt generation-based models to the task of event argument extraction, BART-Gen [28] investigated generative document-level event argument extraction. However, these methods only use naive natural language prompts and have not considered the incorporation of semantic information into the natural language prompts.
To better incorporate label semantics into the model, DEGREE [13] proposed a template-based model that incorporates label semantics into the designed prompts to extract event information. More recently, researchers have started to advance this line of work. For example, GTEE [14] used prefixes [16], which are sequences of tunable continuous prompts, to incorporate contextual type information into the generation-based model. AMPERE [15] proposed to equip the generative model with AMR-aware prefixes to embed the AMR graph, providing more contextual information. To the best of our knowledge, there is no work done on incorporating structural features of complex events into the generative models for BEE. Therefore, in this paper, we propose a novel generative model called GenBEE that integrates complex event structural information for BEE.
@@ III-A Event Prompt Construction
TABLE I: Input event prompt examples for the GE11 dataset.
@@
Most previous works [14, 15] proposed to use manually designed templates for event extraction. However, it is challenging and time-consuming to define event templates for each event type from the biomedical event datasets, which requires domain knowledge. In this work, we propose to employ the Large Language Model (LLM) GPT-4 [23] with designed prompts to generate the event templates for the event types pre-defined in the biomedical event datasets. Figure 3 shows an input prompt that we have designed for generating the event template for the Binding event using GPT-4. The input prompt consists of an “Instruction” that specifies the generation objective for the biomedical event extraction template and a “Basic Template” that outlines the target trigger and argument roles, constructed based on the pre-defined roles given in the technical report associated with the biomedical event dataset. The output of GPT-4 is an updated event template for the corresponding event type, which contains richer information than the basic template for generative event extraction.
Figure 3: Prompts for eliciting knowledge from LLM to construct the event template for the “Binding” event in the GE11 dataset.
@@
where “;” denotes the sequence concatenation, and $[SEP]$ is the special separator token in the PLM (e.g., BART). For each subtask based on $e_{i}$, we construct the input tokens for all the contexts in the dataset.
@@ III-B Structural Prefix Learning
In biomedical texts, events are often complex and interrelated. Identifying the latent relationships among entities across different events could provide important clues for capturing complex event structures, thereby enhancing representation learning. Table II shows the structural prompts [29] constructed in the proposed model to capture different event structures for the Binding event. The structural prompts aim to provide descriptions of different event structures and then elicit the PLM (e.g., BioBERT) to identify complex event structures by capturing the latent relationships among entities across different biomedical events. As shown in Table II, the structural prompts are the same for all event types, except the corresponding event type name specified between the two tags of <T>. In this work, structural prompts are categorized into three groups: General Events, Overlapping Events, and Nested Events. The General Events describe the structural information of common events, guiding the PLM to identify potential interactions between frequently co-occurring events. The Overlapping Events describe the structural information of overlapping events, instructing the PLM to identify trigger words that commonly trigger overlapping events. The Nested Events describe the structural information of nested events, guiding the PLM to identify entities that frequently serve as argument roles in the events while acting as trigger words in others, or vice versa. Overall, the structural prompts are designed to describe the various structural information of biomedical events for generative event extraction.
@@ III-C Generative Model Training
The proposed GenBEE model employs a generative pre-trained language model BART [12] for encoding and decoding. In this module, the prefix $\mathcal{P}_{e_{i}}$ is integrated into both the encoder and decoder of the model. Specifically, from the prefix for event type $e_{i}$, we obtain the additional key and value matrices as $\mathcal{P}_{e_{i}}=K^{e_{i}}=V^{e_{i}}$, which can further be expressed as $K^{e_{i}}=\{k_{1}^{e_{i}},...,k_{l}^{e_{i}}\}$ and $V^{e_{i}}=\{v_{1}^{e_{i}},...,v_{l}^{e_{i}}\}$. Here, $k_{n}^{e_{i}}$ and $v_{n}^{e_{i}}$ $(n\in[1,l])$ are vectors with the same hidden dimension in the Transformer layer. These additional key and value matrices are concatenated with the original key and value matrices in the self-attention layers of the encoder and cross-attention layers of the decoder. As such, when computing the dot-product attention, the query matrix at each position is influenced by these structure-aware prefixes, which will subsequently influence the weightings of the representations for event generation. Note that the prefix $\mathcal{P}_{e_{i}}$ is learnable according to the querying event type $e_{i}$, thereby guiding the BART PLM to differentiate event types with different structural features.
@@ IV-A Datasets
We have conducted experiments on the following three publicly available benchmark datasets, namely MLEE, GE11, and PHEE, for biomedical event extraction. MLEE [19] has 29 event types and 14 argument roles. We use the train/dev/test split given by the data provider. GE11 [20] has 9 event types and 10 role arguments. We use the train/dev/test split given by the shared task and evaluate the performance based on the development set, as the test set is unannotated and the official tool for evaluation is no longer available. PHEE [30] has 2 event types and 16 role arguments. We use the train/dev/test split provided by TextEE [31]. We preprocess each dataset by using the script of TextEE [31]. Table III presents the statistics of the datasets. For evaluation metrics, we follow the previous work [32, 10, 13] to adopt the following two metrics to evaluate the performance: (1) Trigger Classification (Trg-C) F1-score: A trigger is correctly classified if the predicted span and the predicted event type match the gold ones. (2) Argument Classification (Arg-C) F1 score: An argument is correctly classified if the predicted span, event type, and role type match the gold ones.
@@ IV-D Experimental Results
Table IV shows the Trg-C F1-scores and Arg-C F1-scores based on the MLEE, GE11, and PHEE datasets. “Gen” and “Cls” denote Generation-based models and Classification-based models, respectively. The letter “l” in the column PLM is used to denote the large model. Overall, the GenBEE model achieves new state-of-the-art performance on the MLEE and GE11 datasets. Moreover, on the PHEE dataset, our model achieves the second-highest Arg-C F1-score among all baselines, while its Trg-C F1-score is still competitive with other classification-based models. Additionally, we reckon that the simpler event structures with less structural information in the PHEE datasets are the main reason that affects the performance of the GenBEE model.
Table IV also shows that GenBEE significantly outperforms the state-of-the-art generation-based model, DEGREE, in the experiments based on the MLEE, GE11, and PHEE datasets. More specifically, GenBEE demonstrates an improvement of 3.4%, 3.0%, and 2.2% in Trg-C F1-score, and an improvement of 3.3%, 3.9% and 2.4% in Arg-C F1-score, over DEGREE on the MLEE, GE11, and PHEE datasets, respectively. We attribute GenBEE’s performance improvement to the use of event prompts and structural prefixes, which provide event structural information for biomedical event extraction.

### v3

@@ Abstract
Biomedical Event Extraction (BEE) is a challenging task that involves modeling complex relationships between fine-grained entities in biomedical text. BEE has traditionally been formulated as a classification problem. With the recent technological advancements in large language models (LLMs), generation-based models that cast event extraction as a sequence generation problem have attracted much attention from the NLP research communities. However, current generative models often overlook the importance of cross-instance information from complex event structures such as nested events and overlapping events, which contribute quite significantly in the benchmark datasets. In this paper, we propose an event structure-aware generative model called GenBEE, which can capture complex event structures in biomedical text for biomedical event extraction. In particular, GenBEE constructs event prompts that distill knowledge from LLMs for incorporating both label semantics and argument dependency relationships into the proposed model. In addition, GenBEE also generates prefixes with event structural prompts to incorporate structural features for improving the model’s overall performance. We have evaluated the proposed GenBEE model on three widely used biomedical event extraction benchmark datasets, namely MLEE, GE11, and PHEE. Experimental results show that GenBEE has achieved state-of-the-art performance on the MLEE and GE11 datasets, and achieved competitive results when compared to the state-of-the-art classification-based models on the PHEE dataset.
@@ I Introduction
Biomedical Event Extraction (BEE) is a challenging task that involves identifying molecular events from natural language text, where an event typically includes certain biomedical entities, such as genes, proteins, or cellular components [1]. BEE is typically divided into two subtasks, namely biomedical event trigger detection and biomedical event argument extraction. Biomedical event trigger detection aims to identify trigger words and classify them by event types that represent the presence of biomedical events. On the other hand, biomedical event argument extraction aims to identify the arguments and associate them with specific roles in biomedical events, correlating them with the corresponding event triggers. BEE can provide invaluable information to facilitate the curation of knowledge bases and biomolecular pathways [2].
Traditionally, BEE has been formulated as a classification problem [3, 4]. However, these classification-based methods often struggle with encoding label semantics and other forms of weak supervision information [5, 6, 7]. Many current methods extract events in a pipeline manner, where trigger detection and argument extraction are treated as separate phases, which hinders the learning of shared knowledge between subtasks [8]. Recently, generation-based models have been proposed for event extraction (EE) on general domains [9, 10]. These models typically cast the event extraction task as a sequence generation problem using encoder-decoder based pre-trained language models (PLMs) [11, 12] to output conditional generation sequences. These generation-based models benefit from natural language prompts to learn additional semantic information [13]. Apart from employing natural language prompts to enhance semantic learning from the given context, some recent works such as GTEE [14] and AMPERE [15] have also proposed injecting continuous prompts [16, 17] into their frameworks for better contextual representation learning for event argument extraction.
@@
However, most current generation-based EE models [14, 13, 15] have difficulties when tackling events with complex structures, such as nested events. Compared to flat event structures, texts in the biomedical domain often contain more nested event structures [18]. For example, in benchmark datasets like MLEE [19] and GE11 [20], the nested event structures constitute approximately 25% and 28% of all events in the datasets, respectively. Figure 1 shows an example of two nested events identified from the given text. The trigger word “promoters” of the Positive Regulation event (E2) serves as the argument role Site for the trigger word “bind”, which triggers the Binding event (E1). Furthermore, “TCF-1 alpha” serves as the argument role Theme for both events (E1 and E2). In addition, a trigger word may also trigger different events in a sentence, depending on its event type or arguments. Events with such structures are called overlapping events. Identifying these complex structures is important for BEE, as it can reveal the underlying relationships between events. For example, a Site2 argument in a Binding event is likely to be a trigger for a Positive Regulation event.
For generation-based EE models, an intuitive way to incorporate structural information is through prompt engineering [21] by adding natural language prompts that describe complex structural information as part of the input to the generative language model’s encoder. This helps elicit the rich knowledge distributed in the language model. However, Fei et al. [22] pointed out that the mainstay contextual representations in a generative language model’s encoder tend to weaken other high-level information such as structural information, thereby reducing the effects of such information at the decoder. Moreover, it is also unclear whether the generative language model contains sufficient domain knowledge to represent complex biomedical events. Therefore, it is a challenging task to incorporate structural information effectively into generative BEE models.
To address the above issues, we propose an event structure-aware Generative Biomedical Event Extraction model (GenBEE), which aims to capture complex event structures effectively from biomedical texts. First, GenBEE constructs type description prompts to encapsulate the semantics of type labels. Next, it distills the knowledge from large language models (LLMs), such as GPT-4 [23], by creating event template prompts, which incorporate information such as dependencies among argument roles, to enhance contextual representation learning. Additionally, GenBEE introduces a structural prefix learning module that generates structure-aware prefixes, enriched with structural prompts to infuse the structural features.
@@
  • We propose GenBEE, an event structure-aware generative model for biomedical event extraction. Our proposed end-to-end model can distill knowledge from the LLM to incorporate label semantics, dependency relationships among arguments, and shared knowledge between subtasks.
  • In the GenBEE model, we propose a structural prefix learning module to guide the generation process. This module constructs structure-aware prefixes to link continuous prompts with the representation space of generative models. It is worth noting that this plug-and-play module can be incorporated into any biomedical event extraction framework.
@@ II-A Biomedical Event Extraction
Many current methods for biomedical event extraction adopt a “pipeline” strategy, where trigger detection and argument extraction are treated as separate phases. However, “pipeline” methods may result in propagation of errors. To address this issue, Trieu et al. [8] introduced an end-to-end method called DeepEventMine, which simultaneously identifies triggers and allocates roles to entities, thereby mitigating error propagation. However, DeepEventMine requires the full annotation of all entities, which may not always be accessible in datasets. Ramponi et al. [6] considered biomedical event extraction as a sequence labeling task that involves joint training models for trigger detection and argument extraction via multi-task learning. Aside from these classification-based methods, some recent works have formulated event extraction as a question-answering task [24, 25]. This new paradigm transforms traditional classification methods into multiple questioning rounds, producing a natural language answer about a trigger or an argument in each round. Current QA-based event extraction methods are primarily focused on formulating separate questions for different events and argument types. As such, these QA methods lack the capability on capturing dependency information between arguments, which could be effective for assigning argument types.
@@ II-B Generative Event Extraction
Recently, another emerging line of work that treats event extraction as a conditional generation problem [26, 27, 28, 13] has evolved. The development of generation-based event extraction models begins with the reformulation of event extraction problems as a generation task. For example, TANL [26] treated event extraction as translation tasks with augmented labels by using a variety of brackets and vertical bar symbols for representation, and tuned the model to predict these augmented labels. TempGen [27] proposed using templates for constructing role-filler tasks for entity extraction, and generating outputs that fill role entities into non-natural templated sequences. To adapt generation-based models to the task of event argument extraction, BART-Gen [28] investigated generative document-level event argument extraction. However, these methods only use naive natural language prompts and have not considered the incorporation of semantic information into the natural language prompts.
To better incorporate label semantics into the model, DEGREE [13] proposed a template-based model that incorporates label semantics into the designed prompts to extract event information. More recently, researchers have started to advance this line of work. For example, GTEE [14] used prefixes [16], which are sequences of tunable continuous prompts, to incorporate contextual type information into the generation-based model. AMPERE [15] proposed to equip the generative model with AMR-aware prefixes to embed the AMR graph, thereby providing more contextual information. To the best of our knowledge, there is no work done on integrating event structural features of complex events into generative models for BEE. Therefore, in this paper, we propose a novel generative model called GenBEE that incorporates complex event structural information for BEE.
@@ III-A Event Prompt Construction
TABLE I: Example event prompts for the GE11 dataset.
@@
Most previous works [14, 15] proposed to use manually designed templates for event extraction. However, it is challenging and time-consuming to define event templates for each event type in the biomedical event datasets, as domain knowledge is required. In this work, we propose to employ the Large Language Model (LLM) GPT-4 [23] with designed prompts to generate the event templates for the event types pre-defined in the biomedical event datasets. Figure 3 shows an input prompt that we have designed for generating the event template for the Binding event using GPT-4. The input prompt consists of an “Instruction” that specifies the generation objective for the biomedical event extraction template and a “Basic Template” that outlines the target trigger and argument roles. The basic template is constructed based on the pre-defined roles given in the technical report associated with the biomedical event dataset. The output of GPT-4 is an updated event template for the corresponding event type, which contains richer information than the basic template for generative event extraction.
Figure 3: Prompts for eliciting knowledge from LLM to construct the event template for the Binding event in the GE11 dataset.
@@
where “;” denotes the sequence concatenation, and $[SEP]$ is the special separator token in the Pre-trained Language Model (PLM) BART [12]. For each subtask based on $e_{i}$, we construct the input tokens for all the contexts in the dataset.
@@ III-B Structural Prefix Learning
In biomedical texts, events are often complex and interrelated. Identifying the latent relationships among entities across different events could provide important clues for capturing complex event structures, thereby enhancing representation learning. Table II shows the structural prompts [29] constructed in the proposed model to capture different event structures for the Binding event. The structural prompts aim to provide descriptions of different event structures and then elicit the PLM to identify complex event structures by capturing the latent relationships among entities across different biomedical events. As shown in Table II, the structural prompts are the same for all event types, except the corresponding event type name specified between the two tags of <T>. In this work, structural prompts are categorized into three groups: General Events, Overlapping Events, and Nested Events. The General Events describe the structural information of common events, guiding the PLM to identify potential interactions between frequently co-occurring events. The Overlapping Events describe the structural information of overlapping events, instructing the PLM to identify trigger words that commonly trigger overlapping events. The Nested Events describe the structural information of nested events, guiding the PLM to identify entities that frequently serve as argument roles in the events while acting as trigger words in others, or vice versa. Overall, the structural prompts are designed to describe the various structural information of biomedical events for generative event extraction.
@@ III-C Generative Model Training
The proposed GenBEE model employs a generative pre-trained language model BART for encoding and decoding. In this module, the prefix $\mathcal{P}_{e_{i}}$ is integrated into both the encoder and decoder of the model. Specifically, from the prefix for event type $e_{i}$, we obtain the additional key and value matrices as $\mathcal{P}_{e_{i}}=K^{e_{i}}=V^{e_{i}}$, which can further be expressed as $K^{e_{i}}=\{k_{1}^{e_{i}},...,k_{l}^{e_{i}}\}$ and $V^{e_{i}}=\{v_{1}^{e_{i}},...,v_{l}^{e_{i}}\}$. Here, $k_{n}^{e_{i}}$ and $v_{n}^{e_{i}}$ $(n\in[1,l])$ are vectors with the same hidden dimension in the Transformer layer. These additional key and value matrices are concatenated with the original key and value matrices in the self-attention layers of the encoder and cross-attention layers of the decoder. As such, when computing the dot-product attention, the query matrix at each position is influenced by these structure-aware prefixes, which will subsequently influence the weightings of the representations for event generation. Note that the prefix $\mathcal{P}_{e_{i}}$ is learnable according to the querying event type $e_{i}$, thereby guiding the PLM BART to differentiate event types with different structural features.
@@ IV-A Datasets
We have conducted experiments on the following three publicly available benchmark datasets, namely MLEE, GE11, and PHEE, for biomedical event extraction. MLEE [19] has 29 event types and 14 argument roles. We use the train/dev/test split given by the data provider. GE11 [20] has 9 event types and 10 role arguments. We use the train/dev/test split given by the shared task and evaluate the performance based on the development set, as the test set is unannotated and the official tool for evaluation is no longer available. PHEE [30] has 2 event types and 16 role arguments. We use the train/dev/test split provided by TextEE [31] and preprocess each dataset accordingly. Table III presents the statistics of the datasets. For evaluation metrics, we follow the previous work [32, 10, 13] to adopt the following two metrics to evaluate the performance: (1) Trigger Classification (Trg-C) F1-score: A trigger is correctly classified if the predicted span and the predicted event type match the gold ones. (2) Argument Classification (Arg-C) F1 score: An argument is correctly classified if the predicted span, event type, and role type match the gold ones.
@@ IV-D Experimental Results
Table IV shows the performance results based on the MLEE, GE11, and PHEE datasets. “Gen” and “Cls” denote Generation-based models and Classification-based models, respectively. The letter “l” in the column PLM is used to denote the large model. Overall, the GenBEE model achieves new state-of-the-art performance on the MLEE and GE11 datasets. Moreover, on the PHEE dataset, the GenBEE model achieves the second-highest Arg-C F1-score among all baselines, while its Trg-C F1-score is still competitive with other classification-based models. Additionally, we reckon that the simpler event structures with less structural information in the PHEE datasets are the main reason that affects the performance of the GenBEE model.
Table IV also shows that GenBEE significantly outperforms the state-of-the-art generation-based model, DEGREE, in the experiments based on the MLEE, GE11, and PHEE datasets. More specifically, GenBEE demonstrates an improvement of +3.4%, +3.0%, and +2.2% on Trg-C F1-score, and an improvement of +3.3%, +3.9% and +2.4% on Arg-C F1-score, over DEGREE on the MLEE, GE11, and PHEE datasets, respectively. We attribute GenBEE’s performance improvement to the use of event prompts and structural prefixes, which provide event structural information for biomedical event extraction.